OpenStudy (katecc379):

Given the sequence in the table below, determine the sigma notation of the sum for term 4 through term 15.

OpenStudy (katecc379):

is my answer correct

OpenStudy (lord_box):

Yep

OpenStudy (faiqraees):

Correct

OpenStudy (katecc379):

what about this one @Lord_Box

OpenStudy (faiqraees):

Not quite correct

OpenStudy (katecc379):

can you explain

OpenStudy (katecc379):

do i just add all of the sums plugging in 1 through 7?

OpenStudy (lord_box):

Yes

OpenStudy (lord_box):

\[2 - 4 + 8...\]

OpenStudy (faiqraees):

Or you can use the geometric sum formula \[\large\rm Sum=\frac{a(1-r^n)}{1-r} \]

OpenStudy (katecc379):

86?

OpenStudy (lord_box):

yes

OpenStudy (katecc379):

what is a and r

OpenStudy (katecc379):

and n

OpenStudy (lord_box):

a is the initial term, which was 2. r is the common ratio, (-2), n is the number of times, which was 7.

OpenStudy (katecc379):

i don't know what this is

OpenStudy (legomyego180):

It's a geometric series

OpenStudy (katecc379):

how do i solve

OpenStudy (faiqraees):

First identify if its convergent or divergent

OpenStudy (katecc379):

how do i do that

OpenStudy (legomyego180):

There is a sort of formula. http://study.com/cimages/multimages/16/arithmetic-geometric-series-formulas.jpg

OpenStudy (legomyego180):

You have to look at the number inside of the parentheeses and look to see if it is greater or less than one

OpenStudy (katecc379):

so convergent?

OpenStudy (katecc379):

still confused I'm sorry

OpenStudy (faiqraees):

Can you find out the limit of the above expression?

OpenStudy (katecc379):

infinity??

OpenStudy (faiqraees):

\[\large\rm \lim_{x\rightarrow\infty} {\frac{5^x}{6^x}}\]

OpenStudy (katecc379):

5/6?

OpenStudy (faiqraees):

And according to albert test, it says that if the limit is smaller than 1, then the series converges. So is the series convergent or divergent?

OpenStudy (katecc379):

convergent

OpenStudy (katecc379):

so then the sum?

OpenStudy (faiqraees):

Use the geometric formula above

OpenStudy (katecc379):

what formula

OpenStudy (katecc379):

@FaiqRaees

OpenStudy (faiqraees):

Or you can use the geometric sum formula \[\large\rm Sum=\frac{a(1-r^n)}{1-r}\]

OpenStudy (katecc379):

so what is a r and n

OpenStudy (faiqraees):

a=8 n= (any big value like 100) r=5/6

OpenStudy (katecc379):

so 48?

OpenStudy (katecc379):

also is this correct?

OpenStudy (katecc379):

@FaiqRaees

OpenStudy (katecc379):

@Lord_Box

OpenStudy (faiqraees):

It is convergent, but the sum is wrong. See the first term is 1200 so how could the sum be less than first term?

OpenStudy (lord_box):

Try plugging in 1 to both of the convergent series.

OpenStudy (katecc379):

is it c?

OpenStudy (lord_box):

Yep

OpenStudy (faiqraees):

yep